Explain spin of boson and fermion particles.
1
Expert's answer
2020-10-27T11:12:25-0400

Solution

Spin of boson is integer form

As "s=0, \hbar, 2\hbar, 3\hbar......."

And

spin of ferimons us expressed as half integer

"s=\frac{\hbar}{2}, \frac{3\hbar}{2}, \frac{5\hbar}{2}..."